# Trustworthy Conceptual Explanations for Neural Networks in Robot Decision-Making

### v1

@@ Abstract
Black box neural networks are an indispensable part of modern robots. Nevertheless, deploying such high-stakes systems in real-world scenarios poses significant challenges when the stakeholders, such as engineers and legislative bodies, lack insights into the neural networks’ decision-making process. Presently, explainable AI is primarily tailored to natural language processing and computer vision, falling short in two critical aspects when applied in robots: grounding in decision-making tasks and the ability to assess trustworthiness of their explanations. In this paper, we introduce a trustworthy explainable robotics technique based on human-interpretable, high-level concepts that attribute to the decisions made by the neural network. Our proposed technique provides explanations with associated uncertainty scores by matching neural network’s activations with human-interpretable visualizations. To validate our approach, we conducted a series of experiments with various simulated and real-world robot decision-making models, demonstrating the effectiveness of the proposed approach as a post-hoc, human-friendly robot learning diagnostic tool. Code: https://github.com/aditya-taparia/BaTCAVe
@@ I INTRODUCTION
A significant number of models in robotics research are now equipped with deep neural networks (DNNs), with an increasing trend towards end-to-end models [1, 2]. Nevertheless, only a few DNNs are deployed in real-world robots, and those that are deployed mostly focus on tasks such as object detection rather than decision-making or control. This hesitation to use DNNs in real robots is partly due to the high-stakes nature of robot decision-making, where it is unsafe to deploy systems without a clear understanding of their inner workings. Although we do not understand these black-box DNNs, we cannot simply discard them due to their remarkable performance in certain test cases. Hence, we advocate for developing new methods to explain how they work. To this end, instead of focusing on inherently interpretable white-box or gray-box models, this paper focuses on explaining black-box neural networks in robots post-hoc.
While there are many post-hoc explainable techniques proposed by the machine learning community [3, 4, 5], most do not focus on decision-making of a robot or a physical system. For robot decision-making, we want to explain how certain aspects of the input contribute to a particular action or a set of actions. Such explanations help engineers with debugging and legislative bodies with certifying these models. Therefore, to make explanations human-centric, we consider concepts—defined as high-level attributes that help humans understand these black boxes [6]. As an example, the concept of stripes explains why a DNN would classify an image as a zebra. Concepts do not necessarily need to be pixel-level geometric patterns, as in feature attribution methods [3, 7, 4, 8, 5]. For instance, in our experiments, we will show how the concept of darkness of an object can be used to explain collision avoidance decisions of a mobile robot.
If an explainable AI technique provides an explanation about why a robot learning algorithm took a particular decision, why should human trust that explanation? Not only some explanations can be wrong but also there can be multiple explanations for the same decision. Since improving the trustworthiness of a robot explainer is crucial, we propose a method named, Bayesian Testing with Concept Activation Vectors (BaTCAVe), that assigns a score and an associated uncertainty for concepts of interest. The score indicates how well the concept explains the decision and the uncertainty indicates how much to trust the concept. To obtain these scores, we consider a posterior distribution over concept activation vectors, which, due to the non-exponential-family likelihood, is approximated using variational inference. Contributions of the paper are as follows:
@@
    Proposing a theoretical framework to evaluate the trustworthiness of explanations.
@@
    Discovering explainable concepts for diverse robotics decision-making tasks.
Given that failures are inevitable despite efforts to create robust models [9], we believe our work in developing explainability tools will assist roboticists in iteratively improving models to enhance their robustness and safety.
Figure 1: Proposed explainability pipeline. a) The user obtains a score with uncertainty for each “concept.” b) The user provides input concepts $C_{I}$ and test data $\mathbf{x}$. The policy network $f(\cdot)$ returns a score reflecting the alignment of the concept with the network and its uncertainty. Action concepts $C_{A}$ define a subset of the output space for focused analysis, filtering actions based on user interest.
@@ II RELATED WORK
Explainability in robotics: Since robotics is a high-stakes task with many modules that interact with each other, the majority of real-world robots are designed to be interpretable by construction. For instance, they are typically equipped with white-box planning [10] and control [11] algorithms. Even when models are data-driven, they tend to be constructed as gray-boxes—parameters of an inherently interpretable model are estimated using data [12, 13, 14]. As an example, by using a decision tree as the policy, a reinforcement learning algorithm can be made interpretable [12]. Unlike these methods, our focus is developing techniques to probe and explain inherently black-box neural networks in robot decision-making, whether they are used in a modular or an end-to-end fashion. Another line of research has explored explainability of robots for their everyday human users [15, 16]. In contrast, our focus is on engineers and legislators who aim to audit robot learning models and require distinct, actionable explanations.
@@
Explainable AI: Unlike inherently interpretable models, the goal of explainable AI (XAI) is to develop methods for explaining black-box models. XAI techniques focus on how to extract explanations as well as how to represent them. Some techniques achieve explainability by testing input components through perturbations [4] or component removal [7], while others use local approximations of the global decision boundary [7]. XAI methods may leverage gradients [5], weights [17], or layer-wide insights [6] to generate explanations. These explanations can be represented by highlighting specific parts of an image [5], assigning importance scores to an image segment or pixel clusters [3, 7], or by analyzing representative samples [18, 6]. In robot decision-making, highlighting certain parts of an image is not useful as it does not truly explain what aspects of the highlighted pixels is indeed important. For instance, if the XAI technique highlights a car, is it the model or the color that is important? Also, XAI methods typically do not provide uncertainty about their explanations, making them difficult to trust in high-stakes applications such as robotics. Considering the importance of trustworthiness in explanations for robots [19], our approach not only provides explanations but also quantifies their uncertainty.
Concepts: In machine learning, concepts are defined as human-interpretable, high-level attributes. For instance, the concept of stripes is important for a classifier to identify a zebra [6, 20]. While concept-based explanations have been applied in domains such as biomedical imaging [21], they are still applied to typical discrete image classification settings. In this paper, we propose a method to provide explanations for decisions and control commands using concepts. While a recent work discusses continuous explanations [22], it focuses on non-binary concepts rather than non-binary outputs, which are what matters in robot control. We apply the method we propose for decision networks, behavioral cloning, and deep reinforcement learning. Further, our method can provide epistemic uncertainty [23] of explanations, indicating how trustworthy an explanation is.
III METHOD
@@
Similar inputs to a neural network result in similar activation patterns at a given layer. By projecting these activations into a human understandable representation, humans can explain how the neural networks work in certain aspects. To make such a projection, concept activation vectors (CAVs) can be used. They measure how sensitive a classifier’s output is to some activations in the direction of a concept of interest [6]. For instance, if the classification of a zebra image is more sensitive to some images of stripes than images of dots (i.e., the direction), then stripes play a more significant role in classifying a zebra. However, this classical notion in computer vision cannot be used in robotic decision-making for several reasons. First, robot actions can contain continuous control signals, for which CAVs are not defined. Second, they do not provide uncertainty estimates of explanations. Since an explainer will always provide an explanation, we do not know whether to trust them or not without uncertainty. To resolve the first issue, we redefine the concept for robotics in the form of input concepts and action concepts, which are subsequently used to resolve the later issue using Bayesian inference, through which we can estimate the uncertainty of the explanations. As illustrated in Fig. 1, we propose a method that assigns a metric to assess how well a given concept explains robot’s decisions. This metric also comes with an estimation of epistemic uncertainty of the explanation. We propose the following definitions before presenting the method.
Action Concepts: Defining action concepts helps us target specific robot behavior for which we seek explanations. Consider a pre-trained neural network model, $f(\cdot)$, that provides outputs, $\mathbf{y}=f(\mathbf{x})$, with $\mathbf{y}\in\mathbb{R}^{D}$, for an input $\mathbf{x}$. An action concept, $C_{A}$, is a set of conditions that defines a subset of the output space a user is interested in analyzing. Formally, $C_{A}=\oplus_{d=1}^{D}\mathbf{y}_{d}$ for $\mathbf{y}_{d}\in\mathbb{R}_{d}\subseteq\mathbb{R}$ under logical operations, $\oplus$, such as conjunction ($\wedge$), disjunction ($\vee$), and negation ($\lnot$). In a manipulation example, a user might be interested in analyzing the behavior, where the end effector moves down-right while the gripper is open. In such a case, by overloading the notation $x,y$ to represent the cardinal direction in the robot’s physical space and opened gripper distance, $d$, the action concepts can be defined as $C_{A}=\left(\frac{\mathrm{d}x}{\mathrm{d}t}>0\right)\wedge\left(\frac{\mathrm{d}z}{\mathrm{d}t}<0\right)\wedge\left(d>0.5\right)$. Action concepts help us isolate the decisions or behaviors that we are interested in explaining.
Input Concepts: An input concept is a high-level, human interpretable attribute of robot inputs that the user believes is important to explain an action concept. These can be textures, colors, sizes, distances, directions, shapes, objects, etc. Input concepts can be defined based on engineers’ intuition, prior knowledge, or the test cases an engineer or a legislative body is interested in. Additionally, since input concepts can be automatically extracted [20, 24, 25, 26] or generated [26] based on inputs, we consider concept discovery as out-of-scope for this paper. While action concepts are defined as rules, an input concept, $C_{I}$, is defined by a collection of representative inputs, $\{\mathbf{x}_{C_{I}}\}_{m=1}^{M}$. For instance, $M$ images of stripes can be used to explain why an autonomous vehicle slowed down near a crosswalk. Activations at any layer can be computed by passing $\mathbf{x}_{C_{I}}$ through the neural network.
III-A Bayesian Testing with CAVs (BaTCAVe)
@@
Given a collection of input concepts for an action concept, we now derive a score to measure how well each input concept explains the action concept. If an autonomous vehicle inadvertently failed to yield to another vehicle (i.e., action concept), we can test whether the color or the type of the other vehicle (i.e., input concepts) influenced the decision. To formally define the score, let us decompose the neural network into two segments, $f_{-l}(\cdot)$ and $f_{+l}(\cdot)$, with one following the other, at the $l^{\text{text}}$ layer as, $\mathbf{y}=f(\mathbf{x})=f_{+l}(f_{-l}(\mathbf{x}))$ for input $\mathbf{x}$. In other words, $f_{-l}(\cdot)$ are the activations of a DNN at the $l^{\mathrm{th}}$ layer, which are then fed to the rest of the network, $f_{+l}(\cdot)$, to obtain the output, $\mathbf{y}$.
@@
Since our objective is to build a relationship between the activations and human understandable concept inputs for some behaviors (i.e., action concepts), as shown in the plot of Fig. 1, we work in the space of activations (i.e., the space of $f_{-l}(\cdot)$). The dimensionality of the activation space is equal to the number of nodes in the $l^{\text{th}}$ layer. We can obtain the sensitivity of neural network decisions, conditioned by $C_{A}$, w.r.t. layer activations as $\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}$. If we want to measure this sensitivity along a particular direction in the activation space, we have to compute the directional derivative. If this direction is a random variable, $V$, with its realizations, $\mathbf{v}$, is given by the stochastic directional derivative,
@@
$\begin{split}\int\left(\lim_{h\to 0}\frac{f_{+l}^{C_{A}}(f_{-l}(\mathbf{x}+h\mathbf{v}))-f_{+l}^{C_{A}}(f_{-l}(\mathbf{x}))}{h}\cdot p(\mathbf{v})\right)\,\mathrm{d}\mathbf{v}\\ =\int\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}\cdot\mathbf{v}\cdot p(\mathbf{v})\,\mathrm{d}\mathbf{v}.\end{split}$ (1)
@@
If we build a relationships between $V$ and $C_{I}$, we measure the sensitivity of the neural network decisions to human understandable input concepts. Since $V$ is a random variable, it tells us about the uncertainty of this relationship. To build this relationship, let us estimate $\mathbf{v}$ using an external dataset containing inputs from $C_{I}$. To this end, similar to  [6], we collect a concept dataset: $X^{C_{I}}=\{(\mathbf{x}_{m}^{C_{I}},z^{+})\}_{m=1}^{M}$ with input concepts and a different set of inputs $X^{C_{I}^{-}}=\{(\mathbf{x}_{m}^{C_{I}^{-}},z^{-})\}_{m=1}^{M}$ with the positive and negative classes labelled as $z^{+}$ and $z^{-}$, respectively. The negative input concept, $C_{I}^{-}$, can be just another concept or a random collection of inputs, depending on what the user is interested in comparing. How these sets are selected will be described in the Experiments section. See Fig. . 1 for an illustration.
@@
By computing $f_{-l}(\mathbf{x}_{m}^{C_{I}})$ and $f_{-l}(\mathbf{x}_{m}^{C_{I}^{-}})$, we obtain the activation values of the concept dataset. If we can obtain a linear separation between the two classes, then the concept is more unique in the activation space. However, as we will discuss in Section III-B, separation in a high dimensional activation is subjected to uncertainty, we consider a a probabilistic linear separation (i.e., we can draw many separation lines with different probabilities). By setting this probabilistic linear separation the same as $\mathbf{v}$, we obtain the relationship between $V$ and $C_{I}$. To estimate the vector distribution, $\mathbf{v}$, that separates the two classes in the activation space, we apply the Bayes theorem,
@@
Given the positive and negative labels, the likelihood follows a Bernoulli distribution, $z|\mathbf{v}\sim\mathrm{Bernoulli}$. The prior weight distribution is considered to follow a normal distribution, $\mathbf{v}\sim\mathcal{N}$. However, because of the non-conjugate prior, the posterior distribution is not tractable [23, 27]. Hence, we resort to approximate Bayesian inference. Because the activation space of the neural network can be high dimensional, rather than using an MCMC technique, we use variational inference, where we minimize the KL divergence between the true posterior and an approximate posterior distribution, $q(\mathbf{v})$. However, since the true posterior is not known, instead of minimizing the KL divergence, we maximize an evidence lower bound (ELBO),
@@
On a collection of test inputs, $X^{\text{test}}=\{\mathbf{x}_{p}^{\mathrm{test}}\}_{p=1}^{P}$, with relation to (1, we can now compute a score, $s^{C_{A},C_{I},C_{I}^{-}}$,
@@
where $V^{C_{I}}=\{\mathbf{v}_{r}\}_{r=1}^{R}$ are $R$ samples taken from $q(\mathbf{v})$. The higher the score, the better the concept $C_{I}$ explains the test inputs $X^{\text{test}}$. Similarly, the empirical standard deviation reflects the epistemic nature [23] of explanations—how much the model knows that its explanation can be wrong.
III-B Interpreting the Trustworthiness of Explanations

### v2

@@ Abstract
Black box neural networks are an indispensable part of modern robots. Nevertheless, deploying such high-stakes systems in real-world scenarios poses significant challenges when the stakeholders, such as engineers and legislative bodies, lack insights into the neural networks’ decision-making process. Presently, explainable AI is primarily tailored to natural language processing and computer vision, falling short in two critical aspects when applied in robots: grounding in decision-making tasks and the ability to assess trustworthiness of their explanations. In this paper, we introduce a trustworthy explainable robotics technique based on human-interpretable, high-level concepts that attribute to the decisions made by the neural network. Our proposed technique provides explanations with associated uncertainty scores for the explanation by matching neural network’s activations with human-interpretable visualizations. To validate our approach, we conducted a series of experiments with various simulated and real-world robot decision-making models, demonstrating the effectiveness of the proposed approach as a post-hoc, human-friendly robot diagnostic tool. Code: https://github.com/aditya-taparia/BaTCAVe
@@ I INTRODUCTION
A significant number of models in robotics research are now equipped with deep neural networks (DNNs), with an increasing trend towards end-to-end models. Nevertheless, only a few DNNs are deployed in real-world robots, and those that are deployed mostly focus on tasks such as object detection rather than decision-making or control. This hesitation to use DNNs in real robots is partly due to the high-stakes nature of robot decision-making, where it is unsafe to deploy systems without a clear understanding of their inner workings. Although we do not understand these black-box DNNs, we cannot simply discard them due to their remarkable performance in certain test cases. Hence, we advocate for developing new methods to explain how they work. To this end, instead of focusing on inherently interpretable white-box or gray-box models, this paper focuses on explaining black-box neural networks in robots post-hoc.
While there are many post-hoc explainable techniques proposed by the machine learning community [1, 2, 3], most do not focus on decision-making of a robot or a physical system. For robot decision-making, we want to explain how certain aspects of the input contribute to a particular action or a set of actions. Such explanations help engineers with debugging and legislative bodies with certifying these models. Therefore, to make explanations human-centric, we consider concepts—defined as high-level attributes that help humans understand these black boxes [4]. As an example, the concept of stripes explains why a DNN would classify an image as a zebra. Concepts [4] do not necessarily need to be pixel-level geometric patterns, as in feature attribution methods [1, 2, 3, 5]. For instance, in our experiments, we will show how the concept of darkness of an object can be used to explain collision avoidance decisions of a robot.
Figure 1: In this pick-and-place task, users can request a post-hoc explanation for why the robot succeeded (or failed). BaTCAVe probes the policy network to obtain a ranked list of possible explanations, each with an associated likelihood score. In this example, the object’s redness and cylindrical shape are likely contributors to the robot’s actions. The uncertainty intervals indicate how much the user should trust each explanation. Explanations help with model debugging, auditing for regulatory compliance, and building trust.
If an explainable AI method provides an explanation for why a robot learning algorithm made a particular decision, why should a human trust that explanation? As with any explanation, not only some explanations can be wrong but also there can be multiple explanations for the same decision. Since improving the trustworthiness of a robot explainer is crucial, this paper proposes Bayesian Testing with Concept Activation Vectors (BaTCAVe) to provide explanations accompanied by uncertainty. The key contributions are:
@@
    Proposing a framework to evaluate the trustworthiness of explanations using variational inference [6].
@@
    Discovering explainable concepts for diverse robotics decision-making tasks—including, manipulation, navigation, and end-to-end driving—with images, language inputs, or proprioception data.
Given that failures are inevitable [7, 8] despite efforts to create robust models, we believe our work in developing explainability tools will assist roboticists in iteratively improving models to enhance their robustness and safety.
@@ II RELATED WORK
Explainability in robotics: Since robotics is a high-stakes task with many modules that interact with each other, the majority of real-world robots are designed to be interpretable by construction. For instance, they are typically equipped with white-box planning [9] and control [10] algorithms. Even when models are data-driven, they tend to be constructed as gray-boxes—parameters of an inherently interpretable model are estimated using data [11, 12, 13]. As an example, by using a decision tree as the policy, a reinforcement learning algorithm can be made interpretable [11]. Unlike these methods, our focus is on developing techniques to probe and explain inherently black-box neural networks in robot decision-making, whether they are used in a modular or an end-to-end fashion. Another line of research has explored explainability of robots for their everyday human users [14, 15]. In contrast, our focus is on engineers and legislators who aim to audit robot learning models and require distinct, actionable explanations.
@@
Explainable AI: Unlike inherently interpretable models, the goal of explainable AI (XAI) is to develop methods for explaining black-box models. XAI techniques focus on how to extract explanations as well as how to represent them. Some techniques achieve explainability by testing input components through perturbations [2] or component removal [5], while others use local approximations of the global decision boundary [5]. XAI methods may leverage gradients [3], weights [16], or layer-wide insights [4] to generate explanations. These explanations can be represented by highlighting specific parts of an image [3], assigning importance scores to an image segment or pixel clusters [1, 5], or by analyzing representative samples [17, 4]. In robot decision-making, highlighting certain parts of an image is not useful as it does not truly explain what aspects of the highlighted pixels is indeed important. For instance, if the XAI technique highlights a car, is it the model or the color that is important? Also, XAI methods typically do not provide uncertainty about their explanations, making them difficult to trust in high-stakes applications such as robotics. Considering the importance of trustworthiness in explanations for robots [18], our approach not only provides explanations but also quantifies their epistemic uncertainty [6].
III PRELIMINARIES
Concepts: In machine learning, concepts are defined as human-interpretable, high-level attributes such as redness, stripes, young, etc. For instance, the concept of stripes is important for a classifier to identify a zebra [4, 19]. In a classical sense, there are the implicit features a DNN focuses on.
TCAV: With the key observation of similar inputs to a neural network result in similar activation patterns at a given layer, Testing with Concept activation vectors (TCAV) [4] provided the foundation for concept-based explanations. By projecting these activations into a human understandable representation, humans can explain how the neural networks work in certain aspects. To make such a projection, concept activation vectors (CAVs) can be used. They measure how sensitive a classifier’s output is to some activations in the direction of a concept of interest [4]. For instance, if the classification of a zebra image is more sensitive to some images of stripes than images of dots (i.e., the direction in the projected space), then stripes play a more significant role in classifying a zebra than dots. However, this classical notion in computer vision cannot be directly applied to robotic decision-making for several reasons. First, robot actions often involve continuous control signals, for which CAVs are not defined. Second, traditional approaches do not provide uncertainty estimates for their explanations. Since an explainer will always return an explanation, we need an indication of whether the explanation should be trusted.
IV METHOD
Figure 2: The user specifies the robot behavior of interest for analysis using action concepts, $C_{A}$ (e.g., when the robot is acting near the table). The user also provides input concepts, $C_{I}$, (e.g., potential explanations represented as a dictionary of images or proprioception data), and test data, $\mathbf{x}$ (e.g., images or proprioception data). BaTCAVe measures the similarity of activation strengths and directions between each set of input concepts and test data for the given behavior. To do this, the input concepts are linearly separated in the activation space using a Bayesian classifier, resulting in infinite number of classification boundaries, each with a different probability.
IV-A Defining Concepts for Robotic Decision-Making
Robot’s Action Neural Network: We consider a pre-trained neural network in a robot, $f(\cdot)$, that outputs actions, $\mathbf{y}=f(\mathbf{x})$, with $\mathbf{y}\in\mathbb{R}^{D}$, for an input $\mathbf{x}$. For instance, for a task on a 7-DOF manipulator, $f(\cdot)$ is an end-to-end neural network-based policy with $\mathbf{x}$ input images and $\mathbf{y}\in\mathbb{R}^{7}$ control commands to the joints.
Action Concepts: Since a robot takes a sequence of actions in the time horizon $t=0,1,\dots,T$, the entire robot decision-making process cannot be explained at once. Instead, we focus on specific robot behaviors that the user is interested in explaining (Fig. 2). We define an action concept as a specific target robot behavior for which we seek explanations. For instance, why does the manipulator grasp a cube, or why does the autonomous vehicle turns right at a particular road segment? We want to extract what input cues—visual or language—led to these behaviors. Is it the color, geometry, or something else? Formally, an action concept is defined as,
$C_{A}=\left\{t\mid\bigwedge_{d\in D}\text{rule}(\mathbf{y}_{t}^{(d)}),\quad t=\{0,1,\dots\,T\}\right\},$ (1)
where logical operations, $\text{rule}(\cdot)$, such as conjunction ($\wedge$), disjunction ($\vee$), and negation ($\lnot$), are applied across all dimensions of the action vector, $d\in D$, for the entire action series from 0 to $T$. Essentially, we are extracting segments of all trajectories that exhibit a particular behavior. In a manipulation example, a user might be interested in analyzing the behavior, where the end effector moves down-right while the gripper is open. In such a case, by overloading the notation $x,y$ to represent the cardinal direction in the robot’s physical space and opened gripper distance, $d$, the action concepts can be defined as the time steps that obey $\left(\frac{\mathrm{d}x}{\mathrm{d}t}>0\right)\wedge\left(\frac{\mathrm{d}z}{\mathrm{d}t}<0\right)\wedge\left(d>0.5\right)$. Action concepts help us isolate the behaviors that we are interested in explaining.
Input Concepts: An input concept is a high-level, human interpretable attribute of robot inputs that the user believes is important to explain an action concept. These can be textures, colors, sizes, distances, directions, shapes, objects, etc. Input concepts can be defined based on engineers’ intuition, prior knowledge, or the test cases an engineer or a legislative body is interested in. Additionally, since input concepts can be automatically extracted [19, 20, 21] or generated [22] based on inputs, we consider concept discovery as out-of-scope for this paper. While action concepts are defined as rules, an input concept, $C_{I}$, is defined by a collection of representative inputs, $\{\mathbf{x}_{C_{I}}\}_{m=1}^{M}$. For instance, $M$ images of stripes can be used to explain why an autonomous vehicle slowed down near a crosswalk.
@@ IV-A Defining Concepts for Robotic Decision-Making
IV-B Bayesian Testing with CAVs (BaTCAVe)
@@
Given a collection of input concepts for an action concept, we now derive a score to measure how well each input concept explains the action concept. If an autonomous vehicle inadvertently failed to yield to another vehicle (i.e., action concept), we can test if the color or the type of the other vehicle (i.e., input concepts) influenced the decision. To derive a score for an explanation, as shown in Fig. 2, let us decompose the neural network into two segments, $f_{-l}(\cdot)$ and $f_{+l}(\cdot)$, with one following the other, at the $l^{\text{th}}$ layer as, $\mathbf{y}=f(\mathbf{x})=f_{+l}(f_{-l}(\mathbf{x}))$ for input $\mathbf{x}$. In other words, $f_{-l}(\cdot)$ are the activations of a DNN at the $l^{\mathrm{th}}$ layer, which are then fed to the rest of the network, $f_{+l}(\cdot)$, to obtain output $\mathbf{y}$.
@@
Since our objective is to build a relationship between the activations and human understandable concept inputs for some behaviors (i.e., action concepts), as shown in the plot of Fig. 2, we work in the space of activations (i.e., the space of $f_{-l}(\cdot)$). The dimensionality of the activation space is equal to the number of nodes in the $l^{\text{th}}$ layer. We can obtain the sensitivity of neural network decisions, conditioned by $C_{A}$, w.r.t. layer activations as $\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}$. If we want to measure this sensitivity along a particular direction, $\mathbf{v}$, in the activation space, we have to compute the directional derivative $\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}\cdot\mathbf{v}$. If this direction is a random variable, $V$, with its realizations, $\mathbf{v}$, is given by the stochastic directional derivative,
@@
$\begin{split}\mathbb{E}_{V}\left[\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}\cdot\mathbf{v}\right]=\int\frac{\partial f^{C_{A}}(\mathbf{x})}{\partial f_{-l}(\mathbf{x})}\cdot\mathbf{v}\cdot p(\mathbf{v})\,\mathrm{d}\mathbf{v}.\end{split}$ (2)
@@
If we build a relationships between $V$ and $C_{I}$, we measure the sensitivity of the neural network decisions to human understandable input concepts. Since $V$ is a random variable, it also tells us about the uncertainty of this relationship. To build this relationship, let us estimate $\mathbf{v}$ using an external dataset containing inputs from $C_{I}$. To this end, similar to  [4], we collect a concept dataset: $X^{C_{I}}=\{(\mathbf{x}_{m}^{C_{I}},z^{+})\}_{m=1}^{M}$ with input concepts and a different set of inputs $X^{C_{I}^{-}}=\{(\mathbf{x}_{m}^{C_{I}^{-}},z^{-})\}_{m=1}^{M}$ with positive and negative classes labeled as $z^{+}$ and $z^{-}$, respectively. The negative input concept, $C_{I}^{-}$, can be just another concept or a random collection of inputs, depending on what the user is interested in comparing. How these sets are selected will be described in the Experiments section. See Fig. 2 for an illustration.
@@
By computing $f_{-l}(\mathbf{x}_{m}^{C_{I}})$ and $f_{-l}(\mathbf{x}_{m}^{C_{I}^{-}})$, we obtain the activation values of the concept dataset. If we can obtain a linear separation between the two classes, then the concept is more unique in the activation space. However, as we will discuss in Section IV-C, separation in a high dimensional activation is subjected to uncertainty, we consider a probabilistic linear separation (i.e., we can draw many separation lines with different probabilities). By setting this probabilistic linear separation the same as $\mathbf{v}$, we obtain the relationship between $V$ and $C_{I}$. To estimate the vector distribution, $\mathbf{v}$, that separates the two classes in the activation space, we apply the Bayes theorem,
@@
Given the positive and negative labels, the likelihood follows a Bernoulli distribution, $z|\mathbf{v}\sim\mathrm{Bernoulli}$. The prior weight distribution is considered to follow a normal distribution, $\mathbf{v}\sim\mathcal{N}$. However, because of the non-conjugate prior, the posterior distribution is not tractable [6, 23]. Hence, we resort to approximate Bayesian inference. Because the activation space of the neural network can be high dimensional, rather than using an MCMC technique, we use a much faster and scalable method—variational inference [23]—where we minimize the KL divergence between the true posterior and an approximate posterior distribution, $q(\mathbf{v})$. However, since the true posterior is not known, instead of minimizing the KL divergence, we maximize an evidence lower bound (ELBO),
@@
On a collection of test inputs, $X^{\text{test}}=\{\mathbf{x}_{p}^{\mathrm{test}}\}_{p=1}^{P}$, with relation to (2), we can now compute a score, $s^{C_{A},C_{I},C_{I}^{-}}$,
@@
where $V^{C_{I}}=\{\mathbf{v}_{r}\}_{r=1}^{R}$ are $R$ samples taken from $q(\mathbf{v})$.
Interpreting the score: The higher the score, the better the concept $C_{I}$ explains the test inputs $X^{\text{test}}$ or the robot behavior.
Interpreting the uncertainty: Similarly, the empirical standard deviation reflects the epistemic nature [6] of explanations—how much the model knows that its explanation can be wrong. Let us now discuss different variations of uncertainty. Note that, since we considered the full distribution over parameters, the uncertainty is calibrated by construction (our distributions on parameters are on a generalized linear model of the activation space, not on the neural network) [25].
Figure 4: (a) Shows the distribution of input concepts $C_{I}$ selected by the participants for object avoidance task. (b) and (c) shows the effects of fine-tuning and data augmentation, respectively. The higher the score, the better the explanation is. (d) shows samples of dark and light concepts. (e) shows while common XAI methods such as GradCam [3] and LIME [1] can highlight the orange box, they do not reveal what attributes (i.e., input concepts) of the box contribute to the decision of the DNN, making it harder for the engineers to improve the DNN based on the explanations. In contrast, BaTCAVe provides semantically meaningful explanations. (f) highlights the change in confidence over modifying darkness factor in input with models trained with different data augmentation (C-Modification).
Figure 5: A JetBot rollout. We investigate what attributes (i.e., concepts) of the obstacle influenced the decision to turn.
IV-C Interpreting the Trustworthiness of Explanations